OpenStudy (anonymous):

Is it true that if A is invertible and A^T = −A, then (A^−1)^T = −(A^−1).

terenzreignz (terenzreignz):

Transpose of the inverse is the inverse of the transpose :)

terenzreignz (terenzreignz):

So that means... \[\huge \left(A^{-1}\right)^T=\left(A^T\right)^{-1}\]